OpenStudy (anonymous):

A fair coin is flipped four times. What is the probability of getting heads at least once? Write your answer as a simplified fraction.

OpenStudy (anonymous):

pls use the complement approach

OpenStudy (anonymous):

That's what I would have thought anyway. What's the probability of not getting heads at all?

OpenStudy (anonymous):

ah the probability is getting all tails

OpenStudy (anonymous):

Which is?

OpenStudy (anonymous):

it is 1/2*1/2*1/2*1/2

OpenStudy (anonymous):

1/16

OpenStudy (anonymous):

so whts ur reply?

OpenStudy (anonymous):

Yes, that's correct. Now, there are only two possibilities: Either we get no heads or at all, or we get at least one. So, since the probabilities must add to one, what do we get?

OpenStudy (anonymous):

p(of having atlesat one head)=1-p(of having no head)

OpenStudy (anonymous):

1-1/16=15/16

OpenStudy (anonymous):

there ya go

OpenStudy (anonymous):

welll my maths sucks

OpenStudy (anonymous):

Hahah well you handled that fine

OpenStudy (anonymous):

is there any algorith to solve probability ques

OpenStudy (anonymous):

No, they need to be solved on a case-by-case basis.

Directrix (directrix):

You can use Binomial Probability Distributions to solve the problem. It is an algorithm of sorts. P(at least 1 H) = 1 - P(0H) = 1 - C(4,0) (1/2)^0 (1/2)^4 = 1 - 1 (1/16) = 15/16. Note that C(4,0) is combinatorial notation for 4 choose 0. Check out this video on binomial probabilities: http://www.youtube.com/watch?v=xNLQuuvE9ug

OpenStudy (anonymous):

what is a combinatoral notation?

OpenStudy (anonymous):

got it u are taking about combination notation of nCr.....

OpenStudy (anonymous):

murray what is the other notations for combinations

Directrix (directrix):